A box contains 5 green, 2 yellow, and 3 purple balls. Find the probability of obtaining a green or a yellow or a purple ball in a single draw.
Found 2 solutions by stanbon, mangopeeler07:
Answer by stanbon(75887)

A box contains 5 green, 2 yellow, and 3 purple balls. Find the probability of obtaining a green or a yellow or a purple ball in a single draw.
-------------------
You can't miss. The probability is 10/10 = 1
=================
Cheers,
Stan H.

Answer by mangopeeler07(462)

A box contains 5 green, 2 yellow, and 3 purple balls. Find the probability of obtaining a green or a yellow or a purple ball in a single draw.

There are two answers for this question. For one, there is the smart-alec answer that the probability is 1, because those are the only colors possible, so the ball will be one of those colors.

But if you want a non-smart-alec answer (which I assume you do because you are not like me and you don't want to be lazy), do these steps:

Add the amount of balls. 5+2+3=10

Put the amount of whatever color balls over the amount of total balls.
Green=5/10

Yellow=2/10

Purple=3/10

Reduce them all:
Green=1/2
Yellow=1/5
Purple=3/10

The fractions are the probability.